OpenStudy (anonymous):

Factor the expression completely: 20b-16 Help plz i have no idea what to do

OpenStudy (jango_in_dtown):

20b-16=4.5.b-4.4=4(5b-4)

OpenStudy (anonymous):

so what's the answer

OpenStudy (jango_in_dtown):

4(5b-4)

OpenStudy (anonymous):

okay and how did u get that?

OpenStudy (jango_in_dtown):

4 is a common factor in 20b and 16 so i took out 4 common

OpenStudy (anonymous):

okay thankyou and do u think u can help me with one more

OpenStudy (jango_in_dtown):

ok only one more

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

simplify the expression: 4(3x+2)-2

OpenStudy (jango_in_dtown):

4(3x+2)-2=12x+8-2=12x+6

OpenStudy (anonymous):

but u didn't use the order of operations

OpenStudy (jango_in_dtown):

i did 4(3x+2)-2=4.3.x+4.2-2=12x+8-2=12x+6

OpenStudy (anonymous):

so what is the answer?